solve for x: 10 + 2
--- ---- =4
x+2 x+2

Answer by Nihal@SriLanka(22)

10/x+2 + 2/x+2 = 4
multiplying both sides by x+2 we have
10 + 2 = 4(x+2)
i.e. 12= 4x + 8
i.e. 4 = 4x
i.e. x = 1